OpenStudy (anonymous):

Graph all vertical and horizontal asymptotes of the function f(x)=-4x+13/-2x-7

OpenStudy (anonymous):

\[f(x)=\frac{-4x+13}{-2x-7}=\frac{4x-13}{2x+7}\] Vertical asymptotes occur at x-values that will make \(2x+7=0\). To find horizontal asymptotes, evaluate the limits at infinity: \[\lim_{x\to\pm\infty}f(x)\]